# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v1

@@ 4.1. Containers in Multi-Cloud: Roles and Strategies (RQ1)
Multifaceted Role of Containers in Multi-Cloud Environment: Containers fulfill diverse roles within multi-cloud environment, such as application deployment, service hosting, and resource management. They facilitate deployment of applications across a broad range of systems, from web apps to IoT applications. Furthermore, containers contribute to enhancing security in multi-cloud operations. For researchers, this variety of roles signifies a rich area for inquiry, with opportunities to investigate potential security vulnerabilities, management complexities, and performance overheads. For practitioners, understanding these roles can inform efficient and effective deployment strategies, leading to more robust multi-cloud implementations.
Container Technology Features and their Significance: The unique features of container technology, such as lightweight virtualization, compatibility with orchestration tools, and application in edge computing, offer significant advantages. For researchers, these features present opportunities for further exploration, including how these characteristics (e.g., lightweight virtualization) influence container performance and effectiveness in multi-cloud operations. For practitioners, understanding these features can lead to better utilization of container technology and more efficient multi-cloud deployments.
Implementation Strategies and their Real-World Implications: The strategic implementation of containers significantly affects their functionality in multi-cloud environment. Strategies related to deployment, orchestration, security, performance optimization, cloud service management, and data storage can have significant implications. For researchers, these strategies present an area for exploration, including investigation into their real-world implications, benefits, and limitations. For practitioners, a refined understanding of these strategies can lead to more efficient implementations and operations.
Containerization’s Impact on IoT and Application Development: The incorporation of containers into IoT and app development has transform these fields. Containers offer improved software unit encapsulation, portability, and efficiency, which are particularly important in the distributed and resource-constrained nature of IoT. Researchers can further explore potential challenges in detail such as network complexity, security, and the small footprint requirement of IoT devices. For practitioners, understanding the impact of containerization can lead to the development of more robust, efficient, and secure IoT and app solutions.
@@ 4.2. Patterns and Strategies (RQ2)
Containerized applications in multi-cloud environment requires an understanding of prevailing patterns and strategies. In this context, we discuss the key findings related identified patterns and strategies presented in Table 9.
Diverse Architectural Patterns and Communication Strategies: Our study found that Microservice Architecture and Service-Oriented Architecture (SOA) are the most common used architectural patterns for designing container-based applications in multi-cloud environment. These results align with the existing literature that emphasizes the use of MSA and SOA because of their support for modular development, scalability, and efficient resource use (Balalaie et al., 2016; Dragoni et al., 2017). The other patterns frequently reported are Service Mesh and Service Chaining, which are employed as communication and networking patterns, likely due to their ability to manage complex, distributed service-to-service communications in a scalable way (Zhamak, 2018). The frequent use of Blue-Green Deployment and Object Store Service deployment patterns has also been identified in this study, which may stem from their ability to minimize downtime and ensure high availability—essential factors in a cloud-based setting (Fowler, 2010). Implications: Practitioners can enhance the design, scalability, and efficiency of their applications in a multi-cloud environment by understanding and applying these architectural and communication patterns. Researchers might explore how these patterns evolve with the emergence of new technologies and standards.
@@ 4.4. Automation Challenges and Solution Framework (RQ4 - RQ5)
The Automation Challenge-Solution Framework for Containerized Applications in multi-cloud environment (see Table 12) suggests that multi-cloud management for containerized application is a complex task encompassing several dimensions. These include multi-cloud automation, deployment and scaling, resource management, data and application migration, testing and bench marking, standardization and interoperability, application and service management, and runtime and service discovery. This indicates that addressing one challenge may potentially impact the other areas, necessitating a holistic approach to tackle these issues. In the following we discuss the some of the key observations about this framework.
@@
Importance of Standardization and Interoperability: The framework underscores the crucial role of standardization and interoperability in multi-cloud environment, facilitating seamless integration and management across different cloud platforms. Recent literature validates the importance of standardization and interoperability in multi-cloud environment, emphasizing their role in enhancing the efficiency of cloud services (Alonso et al., 2023a). This agrees with our results, further confirming the critical need for these factors in the domain. The need for standardization and interoperability reiterates the necessity for a unified framework or protocol across different cloud environments. It suggests that attention to these aspects can lead to more seamless, efficient, and secure management of multi-cloud services. Implications: For researchers, the results stress the need to investigate methods to achieve better standardization and interoperability in multi-cloud environment. For practitioners, the findings suggest the need to adopt standards and focus on interoperability while selecting and implementing cloud services, ensuring more efficient management across different platforms.
@@ 4.5. Security Challenges and Solutions Framework (RQ4 - RQ5)
The framework proposed in our study provides a comprehensive and systematic approach to addressing the complex security issues associated with containerized applications in multi-cloud environment. In the following, we discuss the key takeaway from security challenges and solution framwrork presented in Section 3.5 and Table 11.
@@ 4.6. Deployment Challenges and Solutions Framework (RQ4 - RQ5)
Table 14 offers a comprehensive overview of challenges and their corresponding solutions encountered during the deployment of containerized applications in multi-cloud environment. In the following, we provide a discussion on key findings pertaining to the challenges and solutions related to the deployment of containerized applications in multi-cloud environment.
@@
Security and Compliance in multi-cloud environment: Security remains paramount in cloud deployments. Fernandes et al. underscore that the labyrinth of diverse platforms and differing standards in multi-cloud configurations amplifies security complexities (Fernandes et al., 2014). Challenges articulated as Security Deployment in Multi-container Environment and Malicious Service Management Deployment in Table 13 accentuate the primacy of security considerations. Propounded solutions like Cross-Container Isolation and Collusion-Resilient Trust Aggregation Technique spotlight the industry’s shift towards advanced and meticulous security protocols within multi-cloud paradigms. The ever-evolving landscape of multi-cloud deployments demands security solutions that are not only rigorous but also malleable and formidable. To attain comprehensive security across diverse platforms, strategies must encapsulate both depth of understanding and breadth of application. Implications: For the academic community, pioneering avant-garde security solutions apt for multi-cloud environment offers a fertile domain of inquiry. Meanwhile, practitioners should enshrine principles of data integrity, trust assurance, and vulnerability mitigation at the heart of their multi-cloud deployment blueprints.
Integration, Monitoring, and Infrastructure Management: As highlighted by Ferrer et al. (Ferrer et al., 2016) the ability to operate seamlessly across disparate cloud providers and maintain flexible infrastructures is paramount in harnessing the full potential of multi-cloud deployments. The presented challenges, including Network Latency Deployment and Vendor Lock-in Prevention Deployment, underscore the indispensability of agility and cross-platform compatibility. Proposed solutions such as VM Selection for Composite Applications and Interoperability Layer Above Cloud Infrastructure accentuate the importance of crafting adaptable infrastructure and formulating strategies to mitigate vendor entrenchment. In the context of multi-cloud environment, flexibility and interoperability transcend mere advantageous attributes; they become pivotal determinants shaping adaptability, scalability, and the sustained success of deployment blueprints. Implications: For the academic community, delving into novel approaches that champion infrastructure pliability and seamless operation across diverse cloud platforms emerges as a prospective research avenue. For industry professionals, anchoring strategies around flexibility and interoperability not only optimizes current deployments but also fortifies them against future technological evolutions and shifting business dynamics.
@@ 4.7. Monitoring Challenges and Solutions (RQ4 - RQ5)
Table 14 presents a comprehensive view of the Monitoring Challenge-Solution Framework for Containerized Applications in multi-cloud environment. It covers a wide range of monitoring challenges within the intricate multi-cloud setting. In the following, we are providing discussion on the key findings.
@@
Strengthening Security, Compliance, and Trust: Ensuring data privacy, continuous security control, and compliance in multi-cloud containerized applications is essential but challenging due to the distributed nature of cloud resources. Existing studies (e.g., (Alonso et al., 2023a) (Raj et al., 2018)) discussed the challenges of ensuring security and compliance across multi-cloud environment. They highlighted the importance of continuous monitoring and enforcement mechanisms to address security and compliance gaps. Solutions presented in our proposed framework, such as using machine learning for performance supervision and employing security assurance platforms, are aligned with existing studies that emphasis on continuous monitoring and enforcement for security and compliance. Implementing advanced techniques like machine learning for monitoring and adopting security assurance platforms can significantly enhance security, compliance, and trust management in multi-cloud containerized applications. Implications: For practitioners, integrate machine learning-based monitoring and adopt security assurance platforms to ensure continuous security control and compliance across diverse cloud environments. For researchers, explore the scalability and effectiveness of machine learning-driven monitoring techniques and investigate novel methods for enforcing security and compliance in multi-cloud setups.
@@ 5.2. External Validity
Concerns about external validity relate to the applicability and generalizability of study findings. Our research provides a thorough examination of containerization in multi-cloud environments, with findings, analyses, and conclusions specifically tailored to this domain. To ensure robust external validity, we developed a study protocol that outlines our research methodology. The literature review spanned a decade, from January 2013 to March 2023, and included peer-reviewed articles from eight preeminent databases in the fields of software engineering and computer science, which are listed in Table 2. While the review is anchored in academic research, which may not encapsulate unpublished industry practices, the systematic approach and comprehensive timeframe of our analysis make the insights valuable for both academia and practitioners. To complement and enrich our findings, future work could include an industrial study (e.g., blogs, white papers) that would provide a broader view of the application of containerization in practice
@@ 6. Related Work
This section reviews the most relevant existing research in terms of secondary studies such as literature surveys, state-of-the-art analysis, systematic reviews, and mapping studies that consolidate published literature on containerization in multi-cloud environments. The review focuses on (i) existing challenges, proposed solutions, and emerging trends detailed in Section 6.1 and (ii) container-based deployment of multi-cloud systems in Section 6.2.
@@ 6.1.1. Challenges, Applications, Tools, and Emerging Trends
Containers in multi-cloud environments provide a standardized and portable way to package, deploy, and run applications, ensuring seamless deployment and management across diverse cloud platforms (Alonso et al., 2023b) (Saxena et al., 2021). One of the earliest SLR-based studies by Pahl et al. (Pahl et al., 2017) on containerized clouds reviewed a total of 46 studies to identify, taxonomically classify and systematically compare the existing research on containers and their application in the context of cloud-based systems. The SLR classified and compared the selected research studies using a conceptual framework to highlight existing trends and needs for future research. SLR results indicate container orchestration, microservice delivery, continuous development and deployment as emerging trends of research on containerized clouds. The SLR in (Alonso et al., 2023b) reviewed a total of 88 studies (published from 2011 - 2021) on multi-cloud systems. Similar survey studies such as (Saxena et al., 2021) concepts, challenges, requirements and future directions for multi-cloud environments are discussed. A survey of existing approaches and solutions provided by different multi-cloud architectures is entailed along with analysis of the pros and cons of different architectures while comparing the same (Imran et al., 2020).
@@ 6.2.1. Orchestration and Deployment
Naweiluo et al. (Zhou et al., 2022) explored containerization in High Performance Computing (HPC) environments, contrasting it with cloud computing. The results of this study indicate that containers enhance application deployment efficiency, but face challenges in HPC due to high security levels and the need for extensive libraries and packages that affect cloud portability, often resulting in vendor lock-ins. The study also provides a survey and taxonomy on containerization and orchestration efforts in HPC, pointing out differences with cloud environments and identifying future research and engineering potentials. A systematic mapping study by Naylor et al.(Bachiega et al., 2018) explores virtualization in container-based cloud deployments. The mapping study, based on a comprehensive review of major databases, identifies a significant research gap in the performance evaluation of containers, underscoring the need for further investigation on cloud deployment. Carlos et al. (Awada, 2018) present an optimization strategy for deploying microservices in multi-cloud environments focusing on minimizing cloud service costs, network latency, and startup time for new microservices. Their approach uses a Non-dominated Sorting Genetic Algorithm II (NSGA-II) compared to a Greedy First-Fit algorithm, demonstrating a 300% improvement with NSGA-II. This highlights its effectiveness in container and VM orchestration, offering a significant enhancement in deployment of multi-cloud solution in containers. Emiliano et al. (Paladi et al., 2018) survey container orchestration, proposing a reference architecture for autonomic orchestrators, and identifying research challenges in the field. Their work emphasizes the importance of container technologies in cloud environments and the need for advanced orchestration solutions to manage complex multi-container applications effectively. Uchechukwu Awada (Guerrero et al., 2018a) reviews container orchestration tools and platforms, comparing the architectures, components, and capabilities of several Container Service Platforms (CSPs) and orchestration tools like Amazon ECS, Kubernetes, Docker Swarm, and Mesos. The study offers insights into the current state of container orchestration and suggests future research directions, serving as a guide for developers and organizations.

### v2

@@ 4.1. Containers in Multi-Cloud: Roles and Strategies (RQ1)
Multifaceted Role of Containers in Multi-Cloud Environment: Containers fulfill diverse roles within multi-cloud environment, such as application deployment, service hosting, and resource management. They facilitate deployment of applications across a broad range of systems, from web apps to IoT applications. Furthermore, containers contribute to enhancing security in multi-cloud operations. Implications: Based on these findings, researchers have opportunities to investigate further areas, such as potential security vulnerabilities, management complexities, and performance overheads. Similarly, practitioners can use this understanding to inform efficient and effective deployment strategies, leading to more robust multi-cloud implementations.
Container Technology Features and their Significance: The unique features of container technology, such as lightweight virtualization, compatibility with orchestration tools, and application in edge computing, offer significant advantages. Implications: Researchers can explore these features further, including how characteristics like lightweight virtualization influence container performance and effectiveness in multi-cloud operations. Practitioners can leverage this understanding for better utilization of container technology and more efficient multi-cloud deployments.
Implementation Strategies and their Real-World Implications: The strategic implementation of containers significantly affects their functionality in multi-cloud environment. Strategies related to deployment, orchestration, security, performance optimization, cloud service management, and data storage can have significant implications. Implications: Researchers may explore these strategies further, including investigations into their real-world implications, benefits, and limitations. Practitioners can refine their understanding of these strategies to achieve more efficient implementations and operations.
Containerization’s Impact on IoT and Application Development: The incorporation of containers into IoT and app development has transform these fields. Containers offer improved software unit encapsulation, portability, and efficiency, which are particularly important in the distributed and resource-constrained nature of IoT. Implications: Researchers can further explore potential challenges in detail such as network complexity, security, and the small footprint requirement of IoT devices. For practitioners, understanding the impact of containerization can lead to the development of more robust, efficient, and secure IoT and app solutions.
@@ 4.2. Patterns and Strategies (RQ2)
Containerized applications in multi-cloud environment requires an understanding of prevailing patterns and strategies. In this context, we discuss the key findings related to the identified patterns and strategies presented in Table 9.
Diverse Architectural Patterns and Communication Strategies: Our study found that Microservice Architecture and Service-Oriented Architecture (SOA) are the most common used architectural patterns for designing container-based applications in multi-cloud environment. These results align with the existing literature that emphasizes the use of MSA and SOA because of their support for modular development, scalability, and efficient resource use (Balalaie et al., 2016; Dragoni et al., 2017). The other patterns frequently reported are Service Mesh and Service Chaining, which are employed as communication and networking patterns, likely due to their ability to manage complex, distributed service-to-service communications in a scalable way (Zhamak, 2018). The frequent use of black-Green Deployment and Object Store Service deployment patterns has also been identified in this study, which may stem from their ability to minimize downtime and ensure high availability—essential factors in a cloud-based setting (fowler2010blackgreen). Implications: Practitioners can enhance the design, scalability, and efficiency of their applications in a multi-cloud environment by understanding and applying these architectural and communication patterns. Researchers might explore how these patterns evolve with the emergence of new technologies and standards.
@@ 4.4. Automation Challenges and Solution Framework (RQ4 - RQ5)
The Automation Challenge-Solution Framework for Containerized Applications in multi-cloud environment (see Table LABEL:tab:Automation-solutions) suggests that multi-cloud management for containerized application is a complex task encompassing several dimensions. These include multi-cloud automation, deployment and scaling, resource management, data and application migration, testing and bench marking, standardization and interoperability, application and service management, and runtime and service discovery. This indicates that addressing one challenge may potentially impact the other areas, necessitating a holistic approach to tackle these issues. In the following we discuss the some of the key observations about this framework.
@@
Importance of Standardization and Interoperability: The framework underscores the crucial role of standardization and interoperability in multi-cloud environment, facilitating seamless integration and management across different cloud platforms. Recent literature validates the importance of standardization and interoperability in multi-cloud environment, emphasizing their role in enhancing the efficiency of cloud services (Alonso et al., 2023b). This agrees with our results, further confirming the critical need for these factors in the domain. The need for standardization and interoperability reiterates the necessity for a unified framework or protocol across different cloud environments. It suggests that attention to these aspects can lead to more seamless, efficient, and secure management of multi-cloud services. Implications: For researchers, the results stress the need to investigate methods to achieve better standardization and interoperability in multi-cloud environment. For practitioners, the findings suggest the need to adopt standards and focus on interoperability while selecting and implementing cloud services, ensuring more efficient management across different platforms.
@@ 4.5. Security Challenges and Solutions Framework (RQ4 - RQ5)
The framework proposed in our study provides a comprehensive and systematic approach to addressing the complex security issues associated with containerized applications in multi-cloud environment. In the following, we discuss the key takeaway from security challenges and solution framework presented in Section 3.5 and Table LABEL:tab:Security-challenges-solutions.
@@ 4.6. Deployment Challenges and Solutions Framework (RQ4 - RQ5)
Table LABEL:tab:monitoring-solutions offers a comprehensive overview of challenges and their corresponding solutions encountered during the deployment of containerized applications in multi-cloud environment. In the following, we provide a discussion on key findings pertaining to the challenges and solutions related to the deployment of containerized applications in multi-cloud environment.
@@
Security and Compliance in multi-cloud environment: Security remains paramount in cloud deployments. Fernandes et al. underscore that the labyrinth of diverse platforms and differing standards in multi-cloud configurations amplifies security complexities (Fernandes et al., 2014). Challenges articulated as Security Deployment in Multi-container Environment and Malicious Service Management Deployment in Table LABEL:tab:Deployment-solutions accentuate the primacy of security considerations. Propounded solutions like Cross-Container Isolation and Collusion-Resilient Trust Aggregation Technique spotlight the industry’s shift towards advanced and meticulous security protocols within multi-cloud paradigms. The ever-evolving landscape of multi-cloud deployments demands security solutions that are not only rigorous but also malleable and formidable. To attain comprehensive security across diverse platforms, strategies must encapsulate both depth of understanding and breadth of application. Implications: For the academic community, pioneering avant-garde security solutions apt for multi-cloud environment offers a fertile domain of inquiry. Meanwhile, practitioners should enshrine principles of data integrity, trust assurance, and vulnerability mitigation at the heart of their multi-cloud deployment blackprints.
Integration, Monitoring, and Infrastructure Management: As highlighted by Ferrer et al. (Ferrer et al., 2016) the ability to operate seamlessly across disparate cloud providers and maintain flexible infrastructures is paramount in harnessing the full potential of multi-cloud deployments. The presented challenges, including Network Latency Deployment and Vendor Lock-in Prevention Deployment, underscore the indispensability of agility and cross-platform compatibility. Proposed solutions such as VM Selection for Composite Applications and Interoperability Layer Above Cloud Infrastructure accentuate the importance of crafting adaptable infrastructure and formulating strategies to mitigate vendor entrenchment. In the context of multi-cloud environment, flexibility and interoperability transcend mere advantageous attributes; they become pivotal determinants shaping adaptability, scalability, and the sustained success of deployment blackprints. Implications: For the academic community, delving into novel approaches that champion infrastructure pliability and seamless operation across diverse cloud platforms emerges as a prospective research avenue. For industry professionals, anchoring strategies around flexibility and interoperability not only optimizes current deployments but also fortifies them against future technological evolutions and shifting business dynamics.
@@ 4.7. Monitoring Challenges and Solutions (RQ4 - RQ5)
Table LABEL:tab:monitoring-solutions presents a comprehensive view of the Monitoring Challenge-Solution Framework for Containerized Applications in multi-cloud environment. It covers a wide range of monitoring challenges within the intricate multi-cloud setting. In the following, we are providing discussion on the key findings.
@@
Strengthening Security, Compliance, and Trust: Ensuring data privacy, continuous security control, and compliance in multi-cloud containerized applications is essential but challenging due to the distributed nature of cloud resources. Existing studies (e.g., (Alonso et al., 2023b) (Raj et al., 2018)) discussed the challenges of ensuring security and compliance across multi-cloud environment. They highlighted the importance of continuous monitoring and enforcement mechanisms to address security and compliance gaps. Solutions presented in our proposed framework, such as using machine learning for performance supervision and employing security assurance platforms, are aligned with existing studies that emphasis on continuous monitoring and enforcement for security and compliance. Implementing advanced techniques like machine learning for monitoring and adopting security assurance platforms can significantly enhance security, compliance, and trust management in multi-cloud containerized applications. Implications: For practitioners, integrate machine learning-based monitoring and adopt security assurance platforms to ensure continuous security control and compliance across diverse cloud environments. For researchers, explore the scalability and effectiveness of machine learning-driven monitoring techniques and investigate novel methods for enforcing security and compliance in multi-cloud setups.
@@ 5.2. External Validity
Concerns about external validity relate to the applicability and generalizability of study findings. Our research provides a thorough examination of containerization in multi-cloud environment, with findings, analyses, and conclusions specifically tailored to this domain. To ensure robust external validity, we developed a study protocol that outlines our research methodology. The literature review spanned a decade, from January 2013 to March 2023, and included peer-reviewed articles from eight preeminent databases in the fields of software engineering and computer science, which are listed in Table 2. While the review is anchored in academic research, which may not encapsulate unpublished industry practices, the systematic approach and comprehensive timeframe of our analysis make the insights valuable for both academia and practitioners. To complement and enrich our findings, future work could include an industrial study (e.g., blogs, white papers) that would provide a broader view of the application of containerization in practice
@@ 6. Related Work
This section reviews the most relevant existing research in terms of secondary studies such as literature surveys, state-of-the-art analysis, systematic reviews, and mapping studies that consolidate published literature on containerization in multi-cloud environment. The review focuses on (i) existing challenges, proposed solutions, and emerging trends detailed in Section 6.1 and (ii) container-based deployment of multi-cloud systems in Section 6.2.
@@ 6.1.1. Challenges, Applications, Tools, and Emerging Trends
Containers in multi-cloud environment provide a standardized and portable way to package, deploy, and run applications, ensuring seamless deployment and management across diverse cloud platforms (Alonso et al., 2023c) (Saxena et al., 2021). One of the earliest SLR-based studies by Pahl et al. (Pahl et al., 2017) on containerized clouds reviewed a total of 46 studies to identify, taxonomically classify and systematically compare the existing research on containers and their application in the context of cloud-based systems. The SLR classified and compared the selected research studies using a conceptual framework to highlight existing trends and needs for future research. SLR results indicate container orchestration, microservice delivery, continuous development and deployment as emerging trends of research on containerized clouds. The SLR in (Alonso et al., 2023c) reviewed a total of 88 studies (published from 2011 - 2021) on multi-cloud systems. Similar survey studies such as (Saxena et al., 2021) concepts, challenges, requirements and future directions for multi-cloud environment are discussed. A survey of existing approaches and solutions provided by different multi-cloud architectures is entailed along with analysis of the pros and cons of different architectures while comparing the same (Imran et al., 2020).
@@ 6.2.1. Orchestration and Deployment
Naweiluo et al. (Zhou et al., 2022) explored containerization in High Performance Computing (HPC) environments, contrasting it with cloud computing. The results of this study indicate that containers enhance application deployment efficiency, but face challenges in HPC due to high security levels and the need for extensive libraries and packages that affect cloud portability, often resulting in vendor lock-ins. The study also provides a survey and taxonomy on containerization and orchestration efforts in HPC, pointing out differences with cloud environments and identifying future research and engineering potentials. A systematic mapping study by Naylor et al.(Bachiega et al., 2018) explores virtualization in container-based cloud deployments. The mapping study, based on a comprehensive review of major databases, identifies a significant research gap in the performance evaluation of containers, underscoring the need for further investigation on cloud deployment. Carlos et al. (Awada, 2018) present an optimization strategy for deploying microservices in multi-cloud environment focusing on minimizing cloud service costs, network latency, and startup time for new microservices. Their approach uses a Non-dominated Sorting Genetic Algorithm II (NSGA-II) compared to a Greedy First-Fit algorithm, demonstrating a 300% improvement with NSGA-II. This highlights its effectiveness in container and VM orchestration, offering a significant enhancement in deployment of multi-cloud solution in containers. Emiliano et al. (Paladi et al., 2018) survey container orchestration, proposing a reference architecture for autonomic orchestrators, and identifying research challenges in the field. Their work emphasizes the importance of container technologies in cloud environments and the need for advanced orchestration solutions to manage complex multi-container applications effectively. Uchechukwu Awada (Guerrero et al., 2018b) reviews container orchestration tools and platforms, comparing the architectures, components, and capabilities of several Container Service Platforms (CSPs) and orchestration tools like Amazon ECS, Kubernetes, Docker Swarm, and Mesos. The study offers insights into the current state of container orchestration and suggests future research directions, serving as a guide for developers and organizations.